OpenStudy (meilendurcer):

Find the area of a parallelogram with base b and height h. b = 82 cm h = 16.6 cm A. 1,361.2 cm^2 B. 6,724 cm^2 C. 137.78 cm^2 D. 98.6 cm^2

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

Don't mean to be rude or anything, but walk her through it, s how her how you did it so she know's how to do it!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The area of all parallelograms is base times height. Usually, the hard part is determining the height, but that was given to you in the problem, so the answer becomes straightforward. So, it becomes just: A = b x h where you have a given base and height.

OpenStudy (anonymous):

A=B * H subsititue in A= 82 * 16.6 A= 1,361.2 cm

OpenStudy (anonymous):

The reason you have b x h can be seen from the diagram where you can fill in the area of the dotted line triangles.

OpenStudy (anonymous):

Actually, this drawing gives a better conceptual rendering of why the area is b x h because you can "shave off" the left-side triangle and place it under the right side so you have a rectangle area of the same: b x h. Hope that helps.